OpenStudy (anonymous):

PLEASEE HELPP!!! A parallel circuit has two 8.0-ohm resistors and a power source of 9.0 volts. If a 12.5-ohm resistor is added to the circuit in parallel, how will the current be affected and what value will it have current will increase 2.97 amps current will increase 0.55 amps current will decrease 2.97 amps cureent will decrease 0.55 amps

OpenStudy (irishboy123):

the 12.5Ohm resistor will have a 9v voltage drop across it. therefore, I = V/R = 9/12.5 = 0.72A will flow through it. total current leaving the battery or cell will be increased to 9/8 + 9/8 + 9/12.5 = 2.97A. i've drawn the circuit just to be clearer.